# Phase-Stable Optical Fiber Links for Quantum Network Protocols

### v1

@@ III.3 Limitations of Phase Stability in Optical Fiber
Due to the finite speed of light in optical fiber, the level of phase noise suppression possible with active feedback is fundamentally limited. A longer fiber leads to a lower feedback bandwidth because the light must travel from the start of the fiber to the end, then back again before a feedback signal is available. Therefore, not only is the speed at which phase fluctuations can be suppressed limited, but the magnitude of the noise suppression is limited across all frequencies. That is, any added fiber length increases the amount of environmental noise coupling to the optical phase and also decreases the ability to suppress the noise. To reduce the phase noise beyond the causal limit, one must break a long fiber into multiple sections, each with its own active feedback stabilization.
@@
In Figure 4 we show the theoretical best phase noise for our deployed fiber (gray trace) based on the measured unlocked phase noise (black trace) and Equation 4. Note that the assumptions of the equation break down at high frequencies, which manifests as a higher than expected phase noise above 10 kHz in the figure.
The experimental data for the locked fiber (purple trace) closely matches the theoretical estimate from about 5 Hz to 1 kHz. Note that the experimental trace dips below the theoretical trace at certain frequencies, which is allowed if the distribution of phase noise is not perfectly uniform along the length of the fiber. At low frequencies the mismatch is most likely due to differential path length changes between the quantum and classical channels and the difference in wavelength. At high frequencies (> 1 kHz) a servo bump is visible that indicates the feedback bandwidth of the fiber stretcher around 10 kHz, which is not present in the theoretical trace.
Figure 4: Phase noise power spectral density of the unlocked 2 km fiber, $S_{\text{unlocked}}(f)$ (black), the theoretical best locked fiber $S_{\text{locked}}(f)$ (gray), and the experimental locked fiber without optical chopping (purple).
@@ IV The quantum channel
To characterize the quantum channel in the network links, we independently stabilize two deployed 2.1 km optical fibers that terminate at the same location (i.e., the same lab). This configuration is used only for link testing; in a functioning quantum network, the fibers would instead terminate at different locations, linking quantum sources. Both fibers are part of a 10-fiber bundle housed in an outdoor-compatible metal jacket, which helps minimize environmentally induced length fluctuations. The bundle is routed through a utility hallway, an environment that introduces more phase noise to the optical signal than a 3.5 km deployed link running across the city of Boulder, CO, making it a good test case for a deployed network.
The same 500 MHz mode-locked laser used to measure the phase noise of the quantum channel (Section III) is attenuated such that there is an average photon number per pulse of << 1, varied from about 0.001 to 0.9 photons per pulse. This ensures that the interference we measure is due to the quantum coherent state and that ultimately a single-photon state will be compatible with our stabilization design. The relative amplitudes going into each fiber are tuned to compensate for different fiber link losses by using a half waveplate and polarizing beam splitter to split the coherent state.
At the other end of the fibers, the quantum and classical channels are split by a 50/50 beam splitter that directs half the amplitude to the final beam splitter where the quantum states from the two network links are combined and interfered. To decrease the channel losses and further isolate the quantum and classical channels, the 50/50 beam splitter may be replaced by a dichroic mirror that splits the 1542 nm and 1550 nm photons. A piezo-driven stage is located in one of the two interferometer arms to slowly change the relative phase of the two quantum channels.
@@ IV.1 Fiber link indistinguishability
The phase delay responsible for sweeping the fringes is driven by a piezo stage, whose displacement has a nonlinear response to voltage. This is the reason the fringes in Fig. 5 have different temporal periods. To correctly fit a sinusoid, it is necessary to add a chirp parameter to the phase, i.e., $f_{fit}(t)=a\sin(b_{3}t^{3}+b_{2}t^{2}+b_{1}t+b_{0})+c$. However, the high quality of the fit in each case indicates that the two optical paths remain highly indistinguishable throughout the measurement despite the fact that neither polarization nor attenuation is actively controlled.
@@ V Isolation of the quantum and classical channels
Achieving such a high degree of isolation is particularly challenging in the case of optical fiber phase stabilization since the quantum and classical light sources must experience the same path noise. The best stabilization would result from a classical light source that has the same polarization, wavelength and bandwidth as the quantum light source, but quantum/classical isolation requires some distinguishable property. Distinguishing the two channels based on polarization is difficult because there is strong coupling between orthogonal polarization states in optical fiber and high isolation would require several stages of polarizers. Isolation based on wavelength [20] is more feasible due to high-quality off-the-shelf dichroic mirrors, but this method is fundamentally limited by the nonlinear scattering of classical photons into the quantum band via the Raman effect [21, 22]. Even with perfect wavelength discrimination, spontaneous Raman scattering would saturate the quantum channel if distinguished between channels by wavelength only. A 2 km silica fiber would scatter 1542 nm light into a 1 nm band around 1550 nm with an efficiency around $10^{-8}$, which would yield an quantum/classical isolation of at most $10^{8}$ in the experiment [23].
@@
The optical filter rejects about a factor of 100 of the classical light at 1542 nm, while transmitting the quantum light at 1550 nm. These wavelengths are close enough that they largely experience the same phase fluctuations in fiber yet far enough that they can be isolated from one another with an optical bandpass filter. Wavelength isolation could be improved by several orders of magnitude with a filter with a steeper band edge. The temporal gating technique relies on the fact that photons from the quantum channel arrive at precisely defined time intervals because they originate from a pulsed laser, whereas the classical photons are distributed across all times since they originate from a CW laser. We choose a gate time of 250 ps, which reduces the classical counts and also the SNSPD background counts by around an order-of-magnitude, as seen in Fig. 7.
Importantly, the second chopper has a duty cycle of 25% compared to 50% for the first chopper. This allows for imperfect chopper synchronization and extra isolation from CW photons that are reflected multiple times through the fiber. However, a lower duty cycle also decreases the quantum channel throughput, reducing the maximum rate of the quantum network. One may decrease the duty cycle of the classical channel chopper while increasing the duty cycle of the quantum channel, though this may also decrease the phase stability of the link.
@@ VI Conclusion and Outlook
VII Disclaimer

### v2

@@ III.3 Limitations of Phase Stability in Optical Fiber
Because of the finite speed of light in optical fiber, the achievable phase-noise suppression in an actively stabilized link is fundamentally limited by the fiber’s round-trip propagation time. As the fiber length increases, the feedback bandwidth decreases, since the system must wait for light to travel to the remote end and back before applying a correction. This delay limits both how rapidly phase fluctuations can be suppressed and the overall level of suppression attainable at all Fourier frequencies. Longer fibers also accumulate proportionally more environmental noise, further reducing stabilization performance. To overcome this causal limit, a long link must be segmented into multiple shorter sections, each equipped with its own active stabilization loop.
@@
In Figure 4 we provides an estimate of the theoretical lower bound on the locked phase noise for our deployed fiber (gray trace) based on the measured unlocked phase noise (black trace) and Equation 4. Note that the assumptions of the equation break down at high frequencies, which manifests as a higher than expected phase noise above 10 kHz in the figure.
The experimental data for the locked fiber (purple trace) closely matches the theoretical estimate from about 5 Hz to 1 kHz. We also observe that the experimental trace dips below the theoretical trace at certain frequencies, which is possible if the phase noise fluctuations are not perfectly uniform along the length of the fiber. At low frequencies the mismatch is most likely due to differential path length changes between the quantum and classical channels as a result of the difference in wavelengths. At high frequencies (> 1 kHz), the observed servo bump indicates that the loop bandwidth of the fiber stretcher is around 10 kHz, which is not present in the theoretical trace.
Figure 4: Phase noise power spectral density of the unlocked 2.1 km fiber, $S_{\text{unlocked}}(f)$ (black), the theoretical best locked fiber $S_{\text{locked}}(f)$ (gray), and the experimental locked fiber without optical chopping (purple).
@@ IV The quantum channel
To characterize the performance of the quantum channel, we independently stabilize two deployed 2.1 km optical fibers that terminate at the same location (i.e., the same lab). This configuration is used only for phase noise measurement; in a functioning quantum network, the fibers would instead terminate at different locations, linking remote quantum sources. In this measurement, both fibers are part of a 10-fiber bundle, housed in an outdoor-compatible metal jacket, which helps minimize environmentally induced length fluctuations. The bundle is routed through a utility hallway, an environment that introduces more phase noise to the optical signal than a 3.5 km deployed link running across the city of Boulder, CO, making it a good test case for a deployed network.
The same 500 MHz mode-locked laser used to measure the phase noise of the quantum channel (Section III) is attenuated such that there is an average photon number per pulse of << 1, varied from about 0.001 to 0.9 photons per pulse. This ensures that the interference fringes we measure are due to the quantum coherent state, and that ultimately a single-photon state will be compatible with our stabilization architecture. The relative amplitudes distributed to each fiber are tuned to compensate for different fiber link losses by using a half waveplate and polarizing beam splitter to split the coherent state.
At the other end of the fibers, the quantum and classical channels are split by a 50/50 beam splitter that directs half the amplitude to the final beam splitter where the quantum states from the two network links are combined and interfered. To decrease the channel losses and further isolate the quantum and classical channels, the 50/50 beam splitter can be replaced by a dichroic mirror that splits the 1542 nm and 1550 nm photons. A piezo-driven stage is located in one of the two interferometer arms to slowly change the relative phase of the two quantum channels.
@@ IV.1 Fiber link indistinguishability
The phase delay responsible for sweeping the fringes is driven by a piezo stage, whose displacement has a nonlinear response to voltage. For this is the reason, the fringes in Fig. 5 have different temporal periods. To correctly fit a sinusoid, it is necessary to add a chirp parameter to the phase, i.e., $f_{fit}(t)=a\sin(b_{3}t^{3}+b_{2}t^{2}+b_{1}t+b_{0})+c$. However, the high quality of the fit in each case indicates that the two optical paths remain highly indistinguishable throughout the measurement despite the fact that neither polarization nor attenuation is actively controlled.
@@ V Isolation of the quantum and classical channels
Achieving such a high degree of isolation is particularly challenging in the case of optical fiber phase stabilization since the quantum and classical light sources must experience the same path noise. The best stabilization would result from a classical light source that has the same polarization, wavelength and bandwidth as the quantum light source, but quantum/classical isolation requires some distinguishable property. The CW stabilization laser is not actively polarization-controlled after propagation through the deployed fiber. Because the stabilization feedback senses path-length fluctuations independent of polarization, this does not impact phase stabilization. A polarizer is placed only in the quantum detection path to project the single-photon states onto a common polarization mode and remove polarization-induced distinguishability. Distinguishing the two channels based on polarization is difficult because there is strong coupling between orthogonal polarization states in optical fiber and high isolation would require several stages of polarizers.
Isolation based on wavelength [36] is more feasible due to high-quality off-the-shelf dichroic mirrors, but this method is fundamentally limited by the nonlinear scattering of classical photons into the quantum band via the Raman effect [27, 32]. Even with perfect wavelength discrimination, spontaneous Raman scattering would saturate the quantum channel if distinguished between channels by wavelength only. A 2.1 km silica fiber would scatter 1542 nm light into a 1 nm band around 1550 nm with an efficiency around $10^{-8}$, which would yield an quantum/classical isolation of at most $10^{8}$ in the experiment [4].
@@
The optical filter rejects about a factor of 100 of the classical light at 1542 nm, while transmitting the quantum light at 1550 nm. These wavelengths are close enough that they largely experience the same phase fluctuations in fiber, yet far enough that they can be isolated from one another with an optical bandpass filter. Wavelength isolation could be improved by several orders of magnitude with a filter with a steeper band edge.
The temporal gating technique relies on the fact that photons from the quantum channel arrive at precisely defined time intervals because they originate from a pulsed laser, whereas the classical photons are distributed across all times since they originate from a CW laser. The temporal gating technique exploits the fact that photons from the quantum channel arrive at well-defined times determined by the pulsed laser repetition rate, whereas classical photons originating from the CW channel are distributed uniformly in time. Here, “gating” refers to a temporal acceptance window applied to time-tagged detection events. We choose a coincidence window of 250 ps, which equates to a duty cycle of 0.125 over the 2 ns modelocked laser pulse spacing. The window is synchronized to the pulsed source, which reduces both classical leakage counts and detector background counts by approximately an order of magnitude, as shown in Fig. 7.
V.1 Chopper duty cycle
Importantly, the chopper before the SNSPD has a duty cycle of 25% compared to 50% for the chopper that chops the CW light. This allows for imperfect chopper synchronization and extra isolation from CW photons that are reflected multiple times through the fiber. However, a lower duty cycle also decreases the quantum channel throughput, reducing the maximum usable rate of the quantum network. One may decrease the duty cycle of the classical channel chopper while increasing the duty cycle of the quantum channel, though this may also decrease the phase stability of the link. This is because a lower duty cycle $D$ reduces the signal-to-noise ratio of the feedback signal and increases detection noise $S_{\phi,det}$ such that $S_{\phi,det}\propto 1/D$. If we assume a simple phase noise model for the fiber $S_{\phi,fiber}\propto 1/f^{2}$, then the frequency $f_{c}$ at which $S_{\phi,det}=S_{\phi,fiber}$ limits the feedback bandwidth. In this case the bandwidth scales with duty cycle as $f_{c}\propto D^{1/2}$, which yields $\Delta\phi_{\mathrm{RMS}}\approx\sqrt{S_{\phi,det}\cdot f_{c}}\propto D^{-1/4}$. If we reduce the CW duty cycle by 10 times from 50% to 5%, this corresponds to an increase in integrated phase noise of about 1.8 times.
[…]
@@ VI Conclusion and Outlook
VII Acknowledgment
The authors acknowledge support from the National Science Foundation QLCI OMA-2016244, National Science Foundation NQVL:QSTD:Pilot 2435378, the National Science Foundation ECCS 2330228, the University of Colorado Quantum Engineering Initiative, and the National Institute of Standards and Technology.
VIII Disclaimer
@@
IX Disclosures
The authors declare no conflicts of interest.
X Data availability
Data underlying the results presented in this paper are not publicly available at this time but may be obtained from the authors upon reasonable request.